Sale of Equipment Equipment was acquired at the beginning of the year at a cost of $32.500. The equipment was depreciated using the double-dedining-balance method based on an estimated usoful lfe of ten years and an estimated residual value of $630. a. What was the depreciation for the first year? b. Assuming the equipment was sold at the end of year 2 for $7.510, determine the gain or loss on the sale of the equipment. c. Journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.
Answer #1

a. Rate of double declining deprDeprecia rate = 10 years = 20%

Depreciation for 1st year = $32500*20% = $6,500

b. Depreciation for year 2 = (32500-6500)*20% = $5200

Net value = $32500 - 6500 - 5200 = $20,800

Loss on sale of equipment = $20800 - $7510 = $13,290

c.

Account Titles Debit Credit
Cash $7,510
Accumulated Depreciation (6500+5200=) $11,700
Loss on sale of equipment $13,290
Equipment $32,500